OpenStudy (anonymous):

Use number 1,2,3,4,5,6,7,8,9 to make 100 only by +,-.division and multiplication.(for fun)

OpenStudy (anonymous):

...

OpenStudy (anonymous):

can we use the same numbers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

we have to use all the numbers, right ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

can we put two numbers together ?

OpenStudy (anonymous):

no you can't :)

OpenStudy (anonymous):

i asked.he said no...

OpenStudy (anonymous):

but you can put all the numbers in a line

OpenStudy (debbieg):

In a line? do you mean that we can use them as digits of larger numbers? Or can we only use them as single-digit numbers?

OpenStudy (anonymous):

1 - 2 + 3 * 4 * 5 + 6 * 7 + 8 - 9 1 - 2 + 12 * 5 + 42 - 1 1 - 2 + 60 + 42 -1 -1 + 102 - 1 -2 + 102 = 100 ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes i think @kelliegirl33 is correct :)

OpenStudy (anonymous):

I used quite a bit of scratch paper for this...lol

OpenStudy (anonymous):

yes you did it and My one is:1+2+3+4+5+6+7+8*9 and I said no to the before of debbie g

OpenStudy (anonymous):

I didn't think of that one.....lol...it is much easier

OpenStudy (anonymous):

=100

OpenStudy (anonymous):

good question Blaze :)

OpenStudy (anonymous):

yes it's correct too :) but i wanted to make 10 by using 5x2 and then try to make it 100 :D

OpenStudy (anonymous):

Thanks found another:1+2*3*4*5/6+7+8*9=100

OpenStudy (anonymous):

1 + 2 * 3 * 4 * 5 / 6 + 7 + 8 * 9 1 + 6 * 4 * 5 / 6 + 7 + 72 1 + 24 * 5 / 6 + 7 + 72 1 + 120/6 + 7 + 72 1 + 20 + 7 + 72 28 + 72 = 100 did I do that right...pemdas and all ?

OpenStudy (anonymous):

superb and thanks